# StreamMOS: Streaming Moving Object Segmentation with Multi-View Perception and Dual-Span Memory

### v1

@@ Abstract
Moving object segmentation based on LiDAR is a crucial and challenging task for autonomous driving and mobile robotics. Most approaches explore spatio-temporal information from LiDAR sequences to predict moving objects in the current frame. However, they often focus on transferring temporal cues in a single inference and regard every prediction as independent of others. This may cause inconsistent segmentation results for the same object in different frames. To overcome this issue, we propose a streaming network with a memory mechanism, called StreamMOS, to build the association of features and predictions among multiple inferences. Specifically, we utilize a short-term memory to convey historical features, which can be regarded as spatial prior of moving objects and adopted to enhance current inference by temporal fusion. Meanwhile, we build a long-term memory to store previous predictions and exploit them to refine the present forecast at voxel and instance levels through voting. Besides, we present multi-view encoder with cascade projection and asymmetric convolution to extract motion feature of objects in different representations. Extensive experiments validate that our algorithm gets competitive performance on SemanticKITTI and Sipailou Campus datasets. Code will be released at https://github.com/NEU-REAL/StreamMOS.git.
@@ I INTRODUCTION
In urban roads, there are often many dynamic objects with variable trajectories, such as vehicles and pedestrians, which create the collision risk for autonomous vehicles. Meanwhile, the moving objects will cause errors in simultaneous localization and mapping (SLAM) [1] as well as bringing challenges for obstacle avoidance [2] and path planning [3]. As a result, online moving object segmentation (MOS) based on LiDAR points has become a crucial task in multiple fields. However, owing to the unordered and sparsity natures of LiDAR points, MOS still faces some challenging cases, especially difficulty in perceiving moving objects with sparse points at a distance.
To tackle the above problem, the mainstream strategy is to exploit spatio-temporal information from LiDAR sequences. For instance, Chen et al. [4] generate residual image in range view (RV), which reflects spatial position of dynamic objects in each frame and can be utilized to perform temporal fusion to predict moving objects. Following the RV-based projection in [4], Sun et al. [5] adopt motion-guided attention to better explore temporal motion cues from residual images. Besides, some works [6, 7] attempt to map point clouds on bird’s eye view (BEV) and ensure consistent object size and movement. Recently, Wang et al. [8] process LiDAR sequences directly via 4D convolution to construct temporal associations while adding instance detection to promote segmentation integrity.
@@
To implement the idea of streaming, we propose a moving object segmentator, called StreamMOS, which encodes object motion cues from multi-view and adopts dual-span memory to transfer historical information. Specifically, different from previous works that map point clouds on one view, we argue that various viewpoints provide more holistic observations of dynamic objects. Thus, we propose a multi-view encoder that applies a cascade structure to iteratively get dense appearance from RV and perceive intuitive motion on BEV, resulting in more distinguishable features of dynamic objects (Fig. 3(b)). Meanwhile, during BEV encoding, we introduce asymmetric convolution with decoupled strategy to better capture vertical and horizontal motion. Then, we use attention mechanism to implement temporal fusion that aligns features from different times and conveys spatial prior to current inference. Besides, due to the inherent uncertainty of neural networks, the output of segmentation decoder may be inconsistent across frames (Fig. 1(c)). To solve this issue, we propose voting mechanism as post-processing to optimize predicted labels. Its core idea is to statistically analyze long-term motion states at the voxel instance levels, and then select the most likely state to update raw point-wise forecasts. In this way, the previous results can be used to refine current predictions, enhancing the temporal continuity and spatial completeness of segmentation together.
In sammary, the contributions of our work are as follows:
@@
  • We propose a multiple projection architecture to capture the object motion and complete appearance from multi-view. We also present a multi-level voting mechanism to refine segmentation results for every voxel and instance.
  • The extensive experiments confirm that our StreamMOS outperforms previous algorithms on the SemanticKITTI (77.8%) and Sipailou Campus (92.5%) while running in real-time. Our code will be available to the community.
@@ II-A Geometric-based Algorithms
The initial LiDAR-based MOS methods could be referred to as the geometric-based approaches, which typically build the map in advance and remove any dynamic objects through estimating occupancy probability and determining visibility. For example, Schauer et al. [10] proposed a ray casting-based approach that counted the hits and misses of scans to update the occupancy situation of the grid map. Afterwards, Pagad et al. [11] constructed an occupancy octree map and proposed a probability update mechanism to obtain clean point clouds by considering the occupancy history. Despite getting promising results, [10, 11] suffer extensive computational burden due to the ray casting and updating voxel one by one. To improve efficiency, several visibility-based [12, 13, 14] algorithms have been developed. Pomerleau et al. [12] identified moving objects by checking whether the points of the pre-built map are occluded by the points in the query frame. Meanwhile, to avoid mismarked ground points as dynamic reported in [12], Kim et al. [13] retained ground points from removed points using a multi-resolution reverting algorithm. Moreover, Lim et al. [15] introduced a visibility-free approach that removed moving traces by computing pseudo occupancy ratio between the query scan and submap in each grid. Although the above methods can distinguish the motion state of objects and clean maps well, they are often performed offline due to requiring a prior map and may be not suitable for real-time applications.
@@ II-B Learning-based Algorithms
Recently, many studies have focused on utilizing learning-based approaches to eliminate dynamic objects online, which only take consecutive frame point clouds as input rather than a pre-built map. Meanwhile, according to data representation, these approaches could be grouped into projection-based and point-based methods. The former converts point clouds into bird’s eye view (BEV) or range view (RV) images, while the latter processes 3D raw points directly.
Specifically, for point-based algorithms, Mersch et al. [9] adopted sparse 4D convolutions to process a series of LiDAR scans and predicted moving objects in each frame. They also employed a binary Bayes filter to fuse multiple predictions in a sliding window. Subsequently, Kreutz et al. [16] proposed an unsupervised approach to address MOS task in stationary LiDAR and viewed it as a multivariate time series clustering problem. Lately, Wang et al. [8] introduced InsMOS to unify detection and segmentation of moving objects into a network, so that the instance cues can be used to improve segmentation integrity. Although they achieved promising performance, the feature extraction of numerous points in [8] may cause high computational costs.
Compared to the mentioned approaches, projection-based algorithms [4, 5, 17, 6, 7] are generally more efficient owing to handling ordered and dense data. For instance, Chen et al. [4] mapped LiDAR scans into spherical coordinates and generated residual images to extract dynamic information in sequence. Sun et al. [5] designed a dual-branch to explore the spatial-temporal information and relieved boundary blurring problem by a point refinement module. Furthermore, Kim et al. [17] achieved higher performance by using extra semantic features. In contrast to range projection, Mohapatra et al. [6] and Zhou et al. [7] utilized BEV projection to obtain a more intuitive motion representation but the serious loss of spatial information still limited their performance. Thus, to address this issue, our StreamMOS exploits a multi-view encoder to capture object motion from BEV and RV in a series manner, which not only allows for complete observation of objects but also alleviates computational effort. Meanwhile, we construct memory banks to pass past knowledge to current inference, resulting in consistent segmentation across a long sequence.
Figure 2: The overall architecture of StreamMOS. (a) Feature encoder adopts a points-wise encoder to extract point features and project them into BEV. Then, the multi-view encoder with cascade structure and asymmetric convolution is employed to extract motion features from different views. (b) Temporal fusion utilizes an attention module to propagate memory feature into the current inference. (c) Segmentation decoder with parameter-free upsampling adopts multi-scale features to predict class labels. (d) Voting mechanism exploits memory predictions to optimize the motion state of each 3D voxel and instance.
@@ III-A Framework Overview
LiDAR-based MOS aims to determine the motion state of each point in the current scan based on the multi-frame point clouds $\{\mathcal{P}_{t-n}\}_{n=0}^{N}$. To this end, existing methods first adopt the relative pose transformations $\{\mathcal{T}_{t-n\rightarrow t}\}_{n=1}^{N}$ provided by the LiDAR odometry to project the history scans $\{\mathcal{P}_{t-n}\}_{n=1}^{N}$ into ego car coordinate system of the current scan $\mathcal{P}_{t}$ and get $\{\mathcal{P}^{\prime}_{t-n}\}_{n=1}^{N}$. Then, they usually feed $\mathcal{P}_{t}$ and $\{\mathcal{P}^{\prime}_{t-n}\}_{n=1}^{N}$ into a network $\Psi$ to fuse spatio-temporal information and predict classification results $\mathcal{M}_{t}\in\mathbb{R}^{V\times 3}$ of all points in $\mathcal{P}_{t}$, where $V\times 3$ refers to the probability that $V$ points belong to three categories, including unknown, static and moving states.
Different from previous approaches that focus on temporal fusion in a single inference, we extra consider the association among multiple inferences and leverage history feature $\mathcal{H}_{t-1}$ and predictions $\{\mathcal{M}_{t-m}\}_{m=1}^{M}$ to raise the quality of current inference. Thus, our method formulates MOS task as follows:
@@
where $N,M$ are the number of historical LiDAR frames and forecasts. Meanwhile, the details of our network are shown in Fig. 2. Specifically, given a series of scans, our StreamMOS first utilizes a multi-view encoder to capture the motion cues from the viewpoints of BEV and RV. Thereafter, we can get a motion feature $\mathcal{F}_{t}$ that reflects spatial information of moving objects in the current frame. Then, we use a temporal fusion module to combine $\mathcal{F}_{t}$ with historical feature $\mathcal{H}_{t-1}$ retained in short-term memory. By doing this, some prior information can be transferred to the current inference and further used to decode movable objects $\mathcal{O}_{t}$ as well as coarse motion state $\mathcal{C}_{t}$ for all points. Finally, we apply a voting mechanism to update $\mathcal{C}_{t}$ with historical results $\{\mathcal{M}_{t-m}\}_{m=1}^{M}$ stored in long-term memory and instance information derived from $\mathcal{O}_{t}$, thereby yielding the refined prediction $\mathcal{M}_{t}$.
@@ III-B1 Preliminaries
Unlike the existing methods that project point clouds into a single view, such as BEV [7] or RV [17], we believe that mapping points to these views simultaneously could capture more complete appearance and obvious motion cues of dynamic objects. Meanwhile, as shown in the bottom of Fig. 2, the points could be considered as the intermediate carrier to transfer information between different perspectives. To achieve these, we use the Point-to-BEV (P2B) and Point-to-Range (P2R) to project point features into 2D plane while using the BEV-to-Point (B2P) and Range-to-Point (R2P) to gather point features from multi-view. Specifically, assuming that the $k^{th}$ 3D point in the $\mathcal{P}_{t}$ is noted as $p_{k}^{3D}=(x_{k},y_{k},z_{k})$, the P2B project it into a rectangular 2D grid and obtain its coordinate $(u_{k}^{b},v_{k}^{b})$ in BEV. For the P2R, the point $p_{k}^{3D}$ with 3D cartesian coordinate is converted into spherical coordinate $p_{k}^{sph}=(r_{k},\theta_{k},\phi_{k})$ and assigned to the 2D grid in RV with coordinate $(u_{k}^{r},v_{k}^{r})$ [18], where $r_{k}$, $\theta_{k}$, $\phi_{k}$ represent distance, zenith and azimuth angle of point $p_{k}^{3D}$. The points falling into the same grid undergo max-pooling to aggregate features. For R2P and B2P, the grid features of RV and BEV are allocated to 3D points using bilinear interpolation within nearby grids.
@@ III-B2 Network Structure
In the feature encoder, we first use a lightweight PointNet [19] as point-wise encoder to process point clouds ($\mathcal{P}_{t},\{\mathcal{P}^{\prime}_{t-n}\}_{n=1}^{N}$) and obtain $\mathcal{E}_{n}\in\mathbb{R}^{V\times C}\ (n\in\{t-N,...,t\})$, where $C$ means the number of channel. Then, for the feature of each frame, we adopt P2B to project them into BEV and concentrate them along the channel dimension to get BEV feature $\mathcal{G}_{t}^{0}\in\mathbb{R}^{W^{b}\times H^{b}\times(N+1)C}$, where $W^{b},H^{b}$ are the predefined width and height of BEV. Afterwards, we feed $\mathcal{G}_{t}^{0}$ into multi-view encoder (MVE) to extract temporal information and capture object motion from different views.
Figure 3: Illustration of Asymmetric Conv Block and multi-view features.
In the lower part of Fig. 2, after downsampling BEV feature $\mathcal{G}_{t}^{l}(l\in\{0,1\})$, we introduce an asymmetric convolution block (ACB) to perceive the movement of objects. As shown in Fig. 3(a), compared to the typical symmetric convolutional kernel (e.g. 3$\times$3), the kernel size of ACB has one side longer (e.g. 3$\times$5 and 5$\times$3). Besides, it decouples feature extraction into the horizontal and vertical directions, defined as follows:
@@
where $f$ and $\odot$ are feature map and concatenation operation. $\text{Conv}_{h}$ and $\text{Conv}_{v}$ mean asymmetric convolutions, which can expand the receptive field and improve perception ability for dynamic objects since they usually have obvious motion in a certain direction. After that, as displayed in Fig. 2, we apply B2P and P2R to project BEV feature $\mathcal{G}^{b}$ into range view and use convolution layers to generate another motion feature $\mathcal{G}^{r}$, which is further remapped into BEV and fuse with $\mathcal{G}^{b}$. Thus, complete motion cues can be encoded by cascade projection.
Thereafter, we stack two MVEs with a BEV-view encoder to get discriminative motion feature $\mathcal{F}_{t}$. Specially, we visualize multi-view features of different MVE layers in Fig. 3(b). It proves that MVE can extract consistent object information across various perspectives, while the deeper layer is capable of suppressing noise and preserving clearer motion features.

### v2

@@ Abstract
Moving object segmentation based on LiDAR is a crucial and challenging task for autonomous driving and mobile robotics. Most approaches explore spatio-temporal information from LiDAR sequences to predict moving objects in the current frame. However, they often focus on transferring temporal cues in a single inference and regard every prediction as independent of others. This may lead to inconsistent segmentation results for the same object across different frames. To solve this issue, we propose a streaming network with a memory mechanism, called StreamMOS, to build the association of features and predictions among multiple inferences. Specifically, we utilize a short-term memory to convey historical features, which can be regarded as spatial priors of moving objects and are used to enhance current inference by temporal fusion. Meanwhile, we build a long-term memory to store previous predictions and exploit them to refine current forecasts at the voxel and instance levels through voting. Besides, we apply multi-view encoder with cascaded projection and asymmetric convolution to extract motion feature of objects in different representations. Extensive experiments validate that our algorithm gets competitive performance on SemanticKITTI and Sipailou Campus datasets.
Index Terms:
Semantic Scene Understanding; Deep Learning Methods; Computer Vision for Transportation
@@ I Introduction
On urban roads, there are often many dynamic objects with variable trajectories, such as vehicles and pedestrians, which create the collision risk for autonomous vehicles. Meanwhile, these moving objects will cause errors in simultaneous localization and mapping (SLAM) [1], as well as pose challenges for obstacle avoidance [2] and path planning [3]. As a result, online moving object segmentation (MOS) based on LiDAR points has become a crucial task in multiple fields. However, owing to the unordered and sparse nature of LiDAR points, MOS still faces some challenges, especially difficulty in perceiving moving objects with sparse points at a distance.
To tackle the above problem, the mainstream strategy is to exploit spatio-temporal information from LiDAR sequences. For instance, Chen et al. [4] generate residual image in range view (RV), which reflects spatial positions of moving objects in each frame and can be utilized to perform temporal fusion to infer motion states of objects. Based on the RV projection in [4], Sun et al. [5] adopt motion-guided attention to better explore temporal motion cues from residual images. Besides, some works [6, 7] attempt to map point clouds to bird’s eye view (BEV) and ensure consistent object size and movement. Recently, Wang et al. [8] process LiDAR sequences directly via 4D convolution to construct temporal associations while adding instance detection to promote segmentation integrity.
@@
To implement the idea of streaming, we propose a moving object segmenter, named StreamMOS, which captures object motion across multiple views and adopts dual-span memory to transfer historical information. Specifically, different from previous works that map point clouds to one view, we argue that various viewpoints provide more holistic observations of dynamic objects. Thus, we propose a multi-view encoder that uses a cascaded structure to iteratively get dense appearance from RV and perceive intuitive motion on BEV, resulting in more discriminative features of moving objects (in Fig. 3(b)). Meanwhile, during BEV encoding, we introduce asymmetric convolution with decoupled strategy to better capture vertical and horizontal motion. Then, we use attention mechanism to implement temporal fusion that aligns features from different times and conveys spatial prior to current inference. Besides, due to the inherent uncertainty of neural networks, the output of segmentation decoder could be inconsistent across frames (Fig. 1(c)). To solve this issue, we propose voting mechanism as post-processing to optimize predicted labels. Its core idea is to statistically analyze long-term motion states at voxel and instance levels, and then select the most likely state to update raw point-wise forecasts. In this way, the previous results can be used to refine current predictions, enhancing the temporal continuity and spatial completeness of segmentation together.
In summary, the contributions of our work are as follows:
@@
  • We propose a multiple projection architecture to capture object motion and complete appearance from different views. We introduce a multi-level voting mechanism to refine segmentation results for every voxel and instance.
  • The experimental results confirm that our StreamMOS outperforms the previous state-of-the-art on the test sets of SemanticKITTI by 1.1% IoU and Sipailou Campus by 1.7% IoU, while achieving competitive running time.
@@ II-A Geometric-based Algorithms
The initial LiDAR-based MOS algorithms can be referred to as the geometric-based approaches, which typically build the map in advance and remove any dynamic objects through estimating occupancy probability and determining visibility. For example, Schauer et al. [10] proposed a ray casting-based approach that counted the hits and misses of scans to update the occupancy situation of the grid map. Afterwards, Pagad et al. [11] constructed an occupancy octree map and proposed a probability update mechanism to obtain clean point clouds by considering the occupancy history. Despite getting promising results, [10, 11] suffer extensive computational burden due to the ray casting and updating voxel one by one. To improve efficiency, several visibility-based [12, 13, 14] algorithms have been developed. Pomerleau et al. [12] identified moving objects by checking whether the points of the pre-built map are occluded by the points in the query frame. Meanwhile, to avoid mismarked ground points as dynamic reported in [12], Kim et al. [13] retained ground points from removed points using a multi-resolution reverting algorithm. Moreover, Lim et al. [15] introduced a visibility-free approach that removed moving traces by computing pseudo occupancy ratio between the query scan and the submap in each grid. Building on [15], Zhang et al. [16] proposed a height coding descriptor, while Lim et al. [17] introduced instance segmentation to maximize the preservation of static points. Although the above methods clean maps well, they are performed offline due to requiring a prior map, making them unsuitable for real-time applications.
@@ II-B Learning-based Algorithms
Recently, many studies have focused on utilizing learning-based approaches to eliminate dynamic objects online, which take only consecutive frame point clouds as input rather than a pre-built map. Meanwhile, according to data representation, these algorithms can be categorized into projection-based and point-based methods. The former converts point clouds into bird’s eye view (BEV) or range view (RV) images, while the latter processes 3D raw points directly.
Specifically, for point-based algorithms, Mersch et al. [9] adopted sparse 4D convolutions to process a series of LiDAR scans and predicted moving objects in each frame. They also employed a binary Bayes filter to fuse multiple predictions in a sliding window. Subsequently, Kreutz et al. [18] proposed an unsupervised approach to address MOS task in stationary LiDAR and viewed it as a multivariate time series clustering problem. Lately, Wang et al. [8] introduced InsMOS to unify detection and segmentation of moving objects into a network, so that the instance cues can be used to improve segmentation integrity. Although they achieved promising performance, the feature extraction of numerous points in [8] may lead to high computational costs.
Compared to the mentioned approaches, projection-based algorithms [4, 5, 19, 6, 7] are generally more efficient owing to handling ordered and dense data. For instance, Chen et al. [4] mapped LiDAR scans into spherical coordinates and generated residual images to extract dynamic information in sequence. Sun et al. [5] used a dual-branch network to encode spatio-temporal information and mitigated boundary blurring problem with a point refinement module. In addition, Kim et al. [19] achieved higher performance by using extra semantic features. In contrast to the RV projection, Mohapatra et al. [6] and Zhou et al. [7] utilized BEV projection to obtain a more intuitive motion representation, but the serious loss of spatial information still limited performance. To solve this issue, our StreamMOS captures object motion from multiple views in a series manner, allowing for complete observation of objects. We also build memory banks to convey historical knowledge, resulting in consistent segmentation across a long sequence.
Figure 2: The overall architecture of StreamMOS. (a) Feature encoder adopts a point-wise encoder to extract point features and project them into BEV. Then, the multi-view encoder with cascaded structure and asymmetric convolution is applied to encode motion features from different views. (b) Temporal fusion utilizes an attention module to propagate memory feature to the current inference. (c) Segmentation decoder with parameter-free upsampling exploits multi-scale features to predict class labels. (d) Voting mechanism leverages memory predictions to optimize the motion state of each 3D voxel and instance.
@@ III-A Framework Overview
LiDAR-based MOS aims to determine the motion state of each point in the current scan based on the multi-frame point clouds $\{\mathcal{P}_{t-n}\}_{n=0}^{N}$. To this end, existing methods first adopt the relative pose transformations $\{\mathcal{T}_{t-n\rightarrow t}\}_{n=1}^{N}$ provided by the LiDAR odometry to project historical scans $\{\mathcal{P}_{t-n}\}_{n=1}^{N}$ into ego car coordinate system of the current scan $\mathcal{P}_{t}$ and get $\{\mathcal{P}^{\prime}_{t-n}\}_{n=1}^{N}$. Then, they usually feed $\mathcal{P}_{t}$ and $\{\mathcal{P}^{\prime}_{t-n}\}_{n=1}^{N}$ into a network $\Psi$ to fuse spatio-temporal information and predict classification results $\mathcal{M}_{t}\in\mathbb{R}^{V\times 3}$ for all points in $\mathcal{P}_{t}$, where $V\times 3$ refers to the probability that $V$ points belong to three categories, including unknown, static and moving states.
Different from previous approaches that focus on temporal fusion in a single inference, we extra consider the association among multiple inferences and apply historical feature $\mathcal{H}_{t-1}$ and predictions $\{\mathcal{M}_{t-m}\}_{m=1}^{M}$ to raise the quality of current inference. Thus, our method formulates MOS task as follows:
@@
where $N,M$ are the number of historical LiDAR frames and forecasts. Meanwhile, the details of our network are shown in Fig. 2. Specifically, given a series of scans, our StreamMOS first leverages the multi-view encoder to capture motion cues from the viewpoints of BEV and RV. Thereafter, we can get a motion feature $\mathcal{F}_{t}$ that reflects spatial information of moving objects in the current frame. Then, we use a temporal fusion module to combine $\mathcal{F}_{t}$ with historical feature $\mathcal{H}_{t-1}$ retained in short-term memory. By doing this, some prior information can be transferred to the current inference and further utilized to decode movable objects $\mathcal{O}_{t}$ and coarse motion state $\mathcal{C}_{t}$ for all points. Finally, we employ a voting mechanism to update $\mathcal{C}_{t}$ using historical results $\{\mathcal{M}_{t-m}\}_{m=1}^{M}$ stored in long-term memory and instance information derived from $\mathcal{O}_{t}$, thereby yielding the refined prediction $\mathcal{M}_{t}$.
@@ III-B1 Preliminaries
Unlike the existing methods that project point clouds into a single view, such as BEV [7] or RV [19], we believe that mapping points to these views simultaneously could capture more complete appearance and obvious motion cues of dynamic objects. Meanwhile, as shown at the bottom of Fig. 2, the points could be considered as the intermediate carrier to transfer information between different perspectives. For this purpose, we adopt Point-to-BEV (P2B) and Point-to-Range (P2R) to map point features into 2D plane, while using BEV-to-Point (B2P) and Range-to-Point (R2P) to collect the point features from multiple planes. To be specific, assuming that the $k^{th}$ 3D point in $\mathcal{P}_{t}$ is denoted as $p_{k}^{3D}=(x_{k},y_{k},z_{k})$, the P2B projects it into a rectangular 2D grid and obtains its coordinate $(u_{k}^{b},v_{k}^{b})$ in BEV. For the P2R, the point $p_{k}^{3D}$ with 3D cartesian coordinate is converted into spherical coordinate $p_{k}^{sph}=(r_{k},\theta_{k},\phi_{k})$ and assigned to the 2D grid in RV with coordinate $(u_{k}^{r},v_{k}^{r})$ [20], where $r_{k}$, $\theta_{k}$, $\phi_{k}$ represent distance, zenith and azimuth angle of point $p_{k}^{3D}$. The points falling into the same grid undergo max-pooling to aggregate features. For R2P and B2P, the grid features of RV and BEV are allocated to 3D points using bilinear interpolation within nearby grids.
@@ III-B2 Network Structure
In the feature encoder, we first use a lightweight PointNet [21] as point-wise encoder to process point clouds ($\mathcal{P}_{t},\{\mathcal{P}^{\prime}_{t-n}\}_{n=1}^{N}$) and obtain $\mathcal{E}_{n}\in\mathbb{R}^{V\times C}\ (n\in\{t-N,...,t\})$, where $C$ means the number of channels. Then, for the feature of each frame, we adopt P2B to project them into BEV and concatenate them along the channel dimension to get BEV feature $\mathcal{G}_{t}^{0}\in\mathbb{R}^{W^{b}\times H^{b}\times(N+1)C}$, where $W^{b},H^{b}$ are the predefined width and height of BEV. Afterwards, we feed $\mathcal{G}_{t}^{0}$ into multi-view encoder (MVE) to extract temporal information and capture object motion from different views.
In the lower part of Fig. 2, after downsampling BEV feature $\mathcal{G}_{t}^{l}(l\in\{0,1\})$, we introduce an asymmetric convolution block (ACB) to perceive the movement of objects. As shown in Fig. 3(a), compared to the typical symmetric convolutional kernel (e.g., 3$\times$3), the kernel size of ACB has one side longer (e.g., 3$\times$5 and 5$\times$3). Besides, it decouples feature extraction into the horizontal and vertical directions, defined as follows:
@@
where $f$ and $\odot$ are feature map and concatenation operation. $\text{Conv}_{h}$ and $\text{Conv}_{v}$ mean asymmetric convolutions, which can expand the receptive field and improve perception ability for moving objects since they usually have distinct motion in a specific direction. After that, as displayed in Fig. 2, we apply B2P and P2R to project BEV feature into the range view and then use convolution layer as range-view encoder to generate another motion feature $\mathcal{G}^{r}$, which is remapped into BEV and combined with $\mathcal{G}^{b}$ along channel dimension. The multi-view feature interaction is then executed in the next encoder layer.
As a result, complete motion information can be extracted through cascaded projection and encoding within two MVEs and an additional BEV encoder, thereby obtaining a discriminative motion feature $\mathcal{F}_{t}$. Specially, we illustrate multi-view features of different MVE layers in Fig. 3(b). It proves that MVE can extract consistent object information across various perspectives, while the deeper layer is capable of suppressing noise and preserving clearer motion features.